What is the percent yield of AlBr3 if 131.6 grams of AlBr3 is produced in when 19.4 g of aluminum reacts? Round to two decimal places and do not include a % sign in your answer.

2Al + 3Br2 → 2AlBr3
1
Expert's answer
2019-03-07T07:11:29-0500

19.4g Xg

2Al + 3Br2 → 2AlBr3

2*27g 2*267g


mtheoretical(AlBr3) = X = 19.4 * 2*267 / (2*27) g = 191.84 g.


The percent yield of AlBr3: η = mpractical(AlBr3) / mtheoretical(AlBr3) * 100% = 131.6 g / 191.84 g * 100% = 68.5973% ≈ 68.60%.


Answer: 68.60.